For which value of k will the roots of 2x^2 + kx + 1 = 0 be real?
Answer by stanbon(48535)

For which value of k will the roots of 2x^2 + kx + 1 = 0 be real?
------------------------
determinant = k^2 - 4*2*1 = k^2-8
-------------------------------
For the roots to be Real the determinant must be non-negative:
k^2-8 >= 0
k^2 >= 8
k >= 2sqrt(2) or k <= -2sqrt(2)
=====================================
Cheers,
Stan H.